What constant angular acceleration is required to reduce an object’s angular velocity from 35.9 rad/s to 11.6 rad/s over 11 revolutions?
Answer #1

Angular acceleration can be calculated from Third equation of rotational kinematics

\omega^2=\omega_0^2 +2\alpha \theta

Theta = 2pi ×11

Omega_0 = 35.9 rad/s

Omega= 11.6 rad/s

So we get,

\alpha=\frac{\omega^2-\omega_0^2 }{2 \theta}

So plug in value we get, angular acceleration

\alpha=-8.35 rad/s^2